question 3 and 4(3) Prove that nen! T($) = lim n70s(s+1)...(s+n) (4) (a) Prove Walliss formula T_2.2 4.4 2m. 2m 2 1.3 3.5 (2m – 1)(2m + 1)
Answer #1


3 we know that r(s) = po fsilent at Re(s) you 1. is called gamma function For n a positive integer and Resyo, we con functionoset-(1144) = et (1-e-t) & et le eteen Plugging this estimate into fg, we deduce that 153 (33) sep (Res +2), which tends to3) x = 10h Sinkeda = k l goth sin k-2 du = H (x-2 This formula is proven using integration by parts. Soth sin kuda = th sinkFor xe [0, %], we know that of sinusl, then Sin 2n+2 na sinantln & Sinann. Thus, 2 Sin 2n+2 dn c 2n+2 m < 1/2 sin anti a / si(25) = (zvy - (алу. (S) (25 +1) ---- (25+ 2n+1) 6 г.(s) ги(st: Гr (25) n“ (n!)ъ. 22 23-1 (2n)! уп с ино (by walles product)